A bullet train leaves a train station and accelerates with 1 g for 60 seconds. Once up to speed it moves with constant speed for another 60 seconds. How far has the train traveled over this trip?

I don’t know what to do when it comes to constant speed!
Answer #1

accelerates with 1 g for 60 seconds

a = 9.8 m/s2

t = 60 sec

s = vot + 1/2at2

s = 0 + (0.5*9.8*602)

d1 = s = 17640 m

v = v0 + at

v = 0 + (9.8*60)

v = 588 m/s

constant speed for another 60 seconds

d2 = v*t

d2 = 588*60

d2 = 35280 m

total trip = 17640 + 35280 = 52920 m

total trip = 52.92 km